OpenStudy (anonymous):

A rational expression has been simplified below. (x+4)(x-2)/9(x-2)= x+4/9 For what values of x are the two expressions equal?

OpenStudy (anonymous):

\[\frac{ (x+4)(x-2) }{9(x-2) }=x+\frac {4}{9}\] The first thing to do here is simplify both sides. X=4/9 is simplified, but you can simplify the other side by canceling out the (x-2) in the numberator and denominator. \[\frac {x+4}{9}=x+\frac{4}{9}\] Now just solve for x. \[9\times(\frac {x+4}{9})= 9 \times (x+\frac{4}{9})\]\[x+4=9x+4\]\[8x=0\] \[x=0\] So when x=0, the two expressions are equal.

OpenStudy (anonymous):

Please verify the problem. The expression should be (x + 4)(x - 2)/9(x - 2)= (x + 4)/9. After simplification by (x - 2) , the remainder is (x + 4)/9. When x = -4, both sides equal to zero When x = 0 , both sides equal to 4/9